OpenStudy (anonymous):

Which ratio is equivalent to ? 9/27

OpenStudy (across):

Perhaps the best way to explain this problem would be to ask you to decompose each number into its primes:\[\frac{9}{27}=\frac{3\cdot3}{3\cdot3\cdot3}=\frac{\not{3}\cdot3}{\not{3}\cdot3\cdot3}=\frac{3}{3\cdot3}=\frac{3}{9}\].